OpenStudy (anonymous):

Solve for X. 9x + 15 = 96

OpenStudy (anonymous):

Someone please help?

OpenStudy (ttp):

x=9 Medal?

OpenStudy (anonymous):

No, x can't equal 9.

OpenStudy (ttp):

And why not?

OpenStudy (anonymous):

Wait what?

OpenStudy (ttp):

Simplifying 9x + 15 = 96 Reorder the terms: 15 + 9x = 96 Solving 15 + 9x = 96 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-15' to each side of the equation. 15 + -15 + 9x = 96 + -15 Combine like terms: 15 + -15 = 0 0 + 9x = 96 + -15 9x = 96 + -15 Combine like terms: 96 + -15 = 81 9x = 81 Divide each side by '9'. x = 9 Simplifying x = 9

OpenStudy (anonymous):

96-15=9x

OpenStudy (anonymous):

And divide by 9.

OpenStudy (anonymous):

Oh lol, thanks. What do you mean by "Medal"?

OpenStudy (anonymous):

That is the simple version.

OpenStudy (anonymous):

I just wanted to solve for X @L-Lawliet-L

OpenStudy (anonymous):

x=9

OpenStudy (ttp):

Mhm by my response is the blue button called "Best response"

OpenStudy (anonymous):

Yes I now know that! Lol.

OpenStudy (anonymous):

I gve you the best resonce.. O.O

OpenStudy (ttp):

Thank you

OpenStudy (anonymous):

She deserves it. @TTP helped you first.

OpenStudy (anonymous):

responce.*

OpenStudy (anonymous):

Your welcome @TTP

OpenStudy (ttp):

:)

OpenStudy (anonymous):

I have another question..

OpenStudy (ttp):

Mhm k ask

OpenStudy (anonymous):

What is it ?

OpenStudy (anonymous):

\[x/7 - 8 = -17\]

OpenStudy (anonymous):

Solve for X

OpenStudy (ttp):

x=17

OpenStudy (anonymous):

And can you walk me through it? Don't just say the answer please?

OpenStudy (anonymous):

Add 8 where?

OpenStudy (ttp):

mhm k you add 8 to both side. Next, Multiply both sides by 7 x=17

OpenStudy (anonymous):

Can you like, write it out?

OpenStudy (anonymous):

And I have another question after thi one.

OpenStudy (anonymous):

You multiply both sides with -1.

OpenStudy (anonymous):

Can you write it out?!?

OpenStudy (anonymous):

\[\frac{ x }{ -1 }=-17\]

OpenStudy (anonymous):

\[x=17\]

OpenStudy (anonymous):

Ugh, this is hard for me to understand.

OpenStudy (ttp):

thanks @L-Lawliet-L

OpenStudy (anonymous):

\[x \over 7 \]

OpenStudy (anonymous):

- 8

OpenStudy (anonymous):

=-17 solve for x

OpenStudy (anonymous):

You are welcome. Your first answer was right. You deserve it.

OpenStudy (anonymous):

Thats the equation

OpenStudy (anonymous):

I didn't thank you, ugh just help me!!

OpenStudy (anonymous):

What does x equal? its not 17!!

OpenStudy (anonymous):

It IS 17.

OpenStudy (anonymous):

No, you're wrong.

OpenStudy (ttp):

And how would you know it's not 17 hun? It is 17

OpenStudy (anonymous):

Prove your answer.

OpenStudy (ttp):

Hold on 1 sec

OpenStudy (anonymous):

If you think it is wrong, why do you ask for our help ? I showed you how to do it.

OpenStudy (anonymous):

Solve for x. x divided by seven minus eight equals negative seventeen

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

cannot be 17, its a multiple choice, and theres no 17

OpenStudy (anonymous):

7-8=-1 You understand this ?

OpenStudy (anonymous):

Omg thats not the equation...

OpenStudy (anonymous):

X can't be negative.

OpenStudy (anonymous):

So you multiply both sides with -1.

OpenStudy (anonymous):

I think i found the misunderstooding.

OpenStudy (ttp):

Sincee it's multiple choice What are your choices?

OpenStudy (anonymous):

\[\frac{ x }{ 7 }-8=-17 \]

OpenStudy (anonymous):

Is it written like this ?

OpenStudy (anonymous):

yes

OpenStudy (ttp):

Is -63 one of your choices?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Add 8 on both sides.

OpenStudy (ttp):

It's -63

OpenStudy (anonymous):

And multiply both sides with 7.

OpenStudy (anonymous):

x=-63

OpenStudy (ttp):

i got my boyfriend to see what his answer was and he say it's -63

OpenStudy (ttp):

So it's -63

OpenStudy (anonymous):

hmm okay, are you sure?

OpenStudy (anonymous):

YES.

OpenStudy (ttp):

Yes i'm sure.

OpenStudy (ttp):

Anymore questions?

OpenStudy (anonymous):

Yeah

OpenStudy (ttp):

Whats the question?

OpenStudy (anonymous):

Jeff bought 3 bags of dog food. He also bought a set of new dog bowls that cost eighteen dollars . He spent a total of forty-five dollars. Write both the verbal and mathematical equation to find out how much Jeff paid for each bag of dog food.

OpenStudy (anonymous):

My answer:The amount of money spent - the cost of the set of new dog bowls = the amount Jeff paid for the dog food bags. 45 - 18= 27/3 27/3= 9 So, Jeff spent $9 dollars for each bag of dog food.

OpenStudy (anonymous):

Is my answer correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay because this is an essay

OpenStudy (ttp):

Mhm okkay, Yes your correct

OpenStudy (anonymous):

I *really* don't wanna fail.

OpenStudy (anonymous):

3*9=27

OpenStudy (ttp):

Is that aall of the queestions?

OpenStudy (anonymous):

Very good.. @L-Lawliet-L

OpenStudy (anonymous):

27+18=45 Do it backwards after it, it does help.

OpenStudy (anonymous):

One more.

OpenStudy (anonymous):

Manuel went shopping and bought 3 pairs of pants. He also spent five dollars on a pair of socks. When he got home, he realized that he had spent a total of seventy-one dollars. What was the cost of each pair of pants?

OpenStudy (ttp):

Glad i could help. You too shoould fan me. Adios!

OpenStudy (anonymous):

71-5=66

OpenStudy (anonymous):

Divide with 3.

OpenStudy (anonymous):

66/3=22

OpenStudy (anonymous):

Thats wutt I have

OpenStudy (anonymous):

71 - 5 / 3 71 - 5 = 66 66/3 = 22. So, Manuel spent $22 dollars on each pair of pants

OpenStudy (anonymous):

Yes, he did.

OpenStudy (anonymous):

ok